How long must a 0.62-mm-diameter aluminum wire be to have a 0.54 A current when connected to the terminals of a 1.5 V flashlight battery?

What is the current if the wire is half this length?

STEP BY STEP
Answer #1

A)

Resistance, R = pL/A

(1.5/0.54) = (2.8 x 10^-8 l) /((pi/4) x 0.00062^2)

Length, l = 29.95 m

B)

Current, I = 1.5/(0.5 x 1.5/0.54) = 1.08 A

Comment in case any doubt please rate my answer....